OpenStudy (anonymous):

What is 5 1/3 · (–4 4/5 )? A. –25 3/5 B. –20 4/15 C. 20 4/15 D. 25 3/5

OpenStudy (anonymous):

Well it is going to be a negative and 5 x 4 is 20 sooo...im guessing b

OpenStudy (anonymous):

One moment

OpenStudy (anonymous):

It's not B

OpenStudy (anonymous):

Nvm just did it on paper its a

OpenStudy (♪chibiterasu):

Can you walk her through it? Don't just give her straight answers.

OpenStudy (anonymous):

5 1/3 X 4 4/5 You convert the whole numbers into fractions when you do that you get 16/3 X 24/5 then you cross multiply and you soon get 16/1 X 8/5 multiply that and you get 128/5then you make the improper fraction proper and you get 25 3/5 Now since a negative times a posotive is a negative you just add the negative sign

OpenStudy (charlotte123):

char to the rescue XD 5 1/3 · (–4 4/5 ) -> 16/3 * -24/5 ---> -384/15 = -25.6 = -25 3/5 ^_^

OpenStudy (anonymous):

Well that was alot shorter version lol

OpenStudy (♪chibiterasu):

Sorry char, Logan gets me medal D:

OpenStudy (anonymous):

When solving this you'd want to change the mixed numbers into improper fractions so for 5 1/3  so 5 1/3 = 16/3 and for -4 4/5  so -4 4/5 = -24/5 Knowing this we can now multiply 16/3 * -24/5 You can cross cancel 3 and -24 because they can both be simplified by divifding by 3 so 3 becomes 1 and -24 becomes -8 16/1 * -8/5 Multiply the numerators and multiply the denominators -128/5 You can divide which would give you -25.6 and .6 of 5 = 3 so 25.6 = 25 3/5, which is A